Solve for X: X + .25X= .4
Answer by The One(11)

X + .25X=0.4


Since X and 0.25X are like terms add 0.25x to x to get 1.25x
1.25x=0.4

Divide each term in the equation by 1.25
1.25x/1.25=0.4/1.25
Cancel the common factor of 1.25
x=0.4/1.25

Divide to perform the operation of division. 0.4 by 1.25 to get 0.32.
x=0.32